NERSCPowering Scientific Discovery Since 1974

March 30

PDSF Users Meeting 3/30/10

Cluster status: Utilization has been fairly light. STAR needs more data.

Outages: Discussed the outage last Wednesday.

Upcoming downtimes: Next Tues memory consumables will be turned on. Website migration will happen in the second week of April.

New hardware: New storage is in for /eliza2; waiting for a couple cables from Dell it will be in production for ATLAS. Discussed trying to coordinate 2 procurements/year.

Retirement plan: All 32-bit nodes are gone now. redhat8 is also gone now.

ATLAS grid activities. Waiting for info from ATLAS.

Other Topics:

- Grid node is ready for STAR

- ALICE vo box is ready

- db nodes for STAR are ready

- pdsfgrid will need to be upgraded at some point